OpenStudy (seattle12345):

Can some one help with this?: The hypotenuse of a right triangle is 50 inches long. One leg is 4 inch(es) longer than the other. Find the lengths of the legs of the triangle. Round your answers to the nearest tenth of an inch. I came up with (x)^2+(x+4)^2=(50)^2 (2x)^2+16=2500 -16 (2x)^2=2484 /2 (x)^2=1242 sqrt x=35.24 and 39.24 But it saying the answer is incorrect, what am I doing wrong?

OpenStudy (waleed_imtiaz):

@seattle12345 what U have written in the second line.....? when U solve te squares of the first line U will get x2+x2+16+8x=2500..... see ur calculation

OpenStudy (anonymous):

See here first line is correct: \[x^2 + (x+4)^2 = 50^2 \] Can you tell me what will you get on expanding this: \[(x + 4)^2 = ??\]

OpenStudy (anonymous):

You have mistaken in expanding this: \[(a + b)^2 = a^2 + b^2 + 2ab\] Use this and expand that..

OpenStudy (anonymous):

@seattle12345 understanding what I am saying or not ??

OpenStudy (anonymous):

No problem you will get a quadratic equation: \[x^2 + x^2 + 16 + 8x = 2500 \implies 2x^2 + 8x - 2500 + 16 = 0\] \[2x^2 + 8x - 2484 = 0 \implies x^2 + 4x - 1242 = 0\]

OpenStudy (anonymous):

You have to go by quadratic formula here: \[D= b^2 - 4ac = 16 + 4968 \implies D = 4984\] \[\sqrt{D} = 70.6\] \[x = \frac{-4 \pm 70.6}{2} \implies x = \frac{-4 + 70.6}{2} \implies x = 33.3\]